# Enhancing the Scalability and Applicability of Kohn-Sham Hamiltonians for Molecular Systems

### v1

@@ 1 Introduction
Consequently, this computational intensity underscores the critical need to develop more efficient methodologies to determining the Hamiltonian without relying on SCF iterations. This challenge has catalyzed interest in leveraging deep learning to predict Hamiltonians directly from atomic configurations while adhering to the inherent symmetries of molecular systems (Unke et al., 2021a; Kochkov et al., 2021; Yu et al., 2023b; Li et al., 2022; Schütt et al., 2019; Gastegger et al., 2020; Hermann et al., 2020; Yin et al., 2024; Zhang et al., 2024). Notably, the Hamiltonian matrix is subject to unitary transformations under molecular rotations due to its spherical harmonics component. To tackle these transformations, researchers have developed SE(3)-equivariant neural networks such as PhisNet (Unke et al., 2021b) and QHNet (Yu et al., 2023b). These networks employ high-order spherical tensors and the Clebsch-Gordon tensor product to construct the predicted Hamiltonians.
@@ 2 Background
For computational efficiency, these orbitals are represented using a basis set $\{\phi_{\alpha}(\mathbf{r})\}_{\alpha=1}^{B}$ depending on the molecular geometry, $B$ denotes the number of basis and varies for different basis sets. The expansion coefficients of the orbitals are organized in a matrix $\mathbf{C}\in\mathbb{R}^{B\times N}$, allowing each orbital $\psi_{i}(\mathbf{r})$ to be expressed as: $\psi_{i}(\mathbf{r})=\sum_{\alpha=1}^{B}C_{\alpha i}\phi_{\alpha}(\mathbf{r}).$ DFT seeks the minimal electronic energy by solving for the optimal coefficient matrix $\mathbf{C}$ via the Kohn-Sham equations, represented as $\mathbf{H}(\mathbf{C})\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon},$ where $\mathbf{H}(\mathbf{C})\in\mathbb{R}^{B\times B}$ denotes the Hamiltonian matrix. Notice that $\mathbf{H}$ is a function of $\mathbf{C}$, and can be computed using density fitting with a complexity of $O(B^{3})$. $\mathbf{S}\in\mathbb{R}^{B\times B}$ is the overlap matrix with elements $S_{\alpha\beta}:=\int\phi^{\dagger}_{\alpha}(\mathbf{r})\phi_{\beta}(\mathbf{r})\mathrm{d}\mathbf{r}$, where $\phi^{\dagger}$ denotes the complex conjugate of $\phi$. $\bm{\epsilon}$ represents a diagonal matrix containing the orbital energies.
This forms a generalized eigenvalue problem, where $\mathbf{C}$ comprises the eigenvectors and the diagonal elements of $\bm{\epsilon}$ are the eigenvalues. However, the solution to this problem is complicated by the interdependence between $\mathbf{H}(\mathbf{C})$ and $\mathbf{C}$. To resolve this, traditional DFT employs the self-consistent field (SCF) method, an iterative process that refines the coefficients $\mathbf{C}^{(k)}$ through successive approximations of the Hamiltonian matrix $\mathbf{H}^{(k)}$. Each iteration commences by computing $\mathbf{H}^{(k)}$ leveraging $\mathbf{C}^{(k-1)}$ and solves a new eigenvalue problem:
@@ 3 Wavefunction-Alignment Loss
Assume $\lambda_{\text{min}}(S)$ of the overlap matrix $S$ scales as: $\lambda_{\text{min}}(S)=c+\frac{A}{1+\left(\dfrac{B}{N_{0}}\right)^{\alpha}},$ where $c>0$, $A>0$, and $\alpha>0$ describe the saturation behavior. Then, for a perturbation matrix $\Delta H$, the eigenvalue perturbation is bounded by:
$\left|\lambda_{i}(\hat{\mathbf{H}},\mathbf{S})-\lambda_{i}(\mathbf{H},\mathbf{S})\right|\leq O\left(\dfrac{1+\left(\dfrac{B}{N_{0}}\right)^{\alpha}}{c\left[1+\left(\dfrac{B}{N_{0}}\right)^{\alpha}\right]+A}\left(\sigma B^{1/2}+B|\mu|\right)\right).$
@@
In light of this, we propose a novel loss function: the Wavefunction Alignment Loss. It is designed to preserve the integrity of the eigenstructure related to molecular orbitals. Let $\hat{\bm{\epsilon}}$ and $\bm{\epsilon}^{\star}$ represent the eigenvalues (orbital energies) of $\hat{\mathbf{H}}$ and $\mathbf{H}^{\star}$, respectively, and $\hat{\mathbf{C}}$ and $\mathbf{C}^{\star}$ denote the corresponding eigenvectors (basis coefficients). We define a primary form of the loss function by directly applying the Frobenius norm to the eigenvalues $\mathcal{L}_{\mathrm{align}}=\frac{1}{n}\sum_{i=1}^{n}\|\hat{\bm{\epsilon}}_{{}_{(i)}}-\bm{\epsilon}_{{}_{(i)}}^{\star}\|^{2}_{\mathrm{F}},$ where $\hat{\bm{\epsilon}}_{{}_{(i)}}$ and $\bm{\epsilon}_{{}_{(i)}}^{\star}$ are derived from solving the generalized eigenvalue problems: $\hat{\mathbf{H}}\hat{\mathbf{C}}=\mathbf{S}\hat{\mathbf{C}}\hat{\bm{\epsilon}}$ and $\mathbf{H}^{\star}\mathbf{C}^{\star}=\mathbf{S}\mathbf{C}^{\star}\bm{\epsilon}^{\star}$, respectively.
@@
When the overlap matrix $\mathbf{S}$ is ill-conditioned, the eigenvalues $\bm{\epsilon}^{\star}$ computed by Algorithm 1 can suffer from significant roundoff errors. To mitigate this, we modify the algorithm by replacing the Cholesky decomposition of $\mathbf{S}$ with its eigen (Schur) decomposition $\mathbf{V}^{\top}\mathbf{S}\mathbf{V}=\mathbf{\Sigma}$, where $\mathbf{V}$ is the matrix of eigenvectors and $\mathbf{\Sigma}$ is the diagonal matrix of eigenvalues. We then substitute $\mathbf{G}$ with $\mathbf{V\Sigma}^{-1/2}$. This modification effectively reorders the entries of $\mathbf{M}^{\star}$, placing larger values towards the upper left-hand corner, thereby enhancing the precision in computing smaller eigenvalues (Golub & Van Loan, 2013; Wilkinson, 1988). Our final algorithm is given in Algorithm 2.
@@
where the subscript $(\cdot)_{{}_{(i)}}$ denotes the $i$-th sample. This modified loss function explicitly penalizes deviations from the expected eigenstructure. It is important to note that while $(\mathbf{C}^{\star})^{\top}\hat{\mathbf{H}}\mathbf{C}^{\star}$ may not yield a diagonal matrix, the diagonal nature of $\bm{\epsilon}^{\star}$ implicitly enforces the predicted eigenstructure through the non-diagonal elements of the loss function.
@@ Appendix B Additional Related Work
Equivariant Graph Neural Networks (EGNNs)  It is often desired that machine learning (ML) models exhibit equivariance to rotations, translations, or reflections, which guarantee that they respect certain physical symmetries. Foundational work introduced group equivariant convolutional neural networks (ECNNs) (Cohen & Welling, 2016), whose importance was underscored by a proof that equivariance and convolutional structure are equivalent given certain ordinary constraints (Kondor & Trivedi, 2018). One appealing way to implement convolution is with geometric graph neural networks (geometric GNNs), which apply naturally to atomic systems by encoding them as graphs embedded in $\mathbb{R}^{3}$. Two important families of geometric GNNs are invariant GNNs and Cartesian equivariant GNNs. Over the past several years, invariant GNNs have achieved state-of-the-art results in predicting properties of molecules, crystals, and other materials  (Schütt et al., 2018) (Sanyal et al., 2018) (Chen et al., 2019) (Gasteiger et al., 2020) (Liu et al., 2022) (Gasteiger et al., 2021) (Wang et al., 2022), as well as in predicting the folding structure of proteins (Jumper et al., 2021). Cartesian equivariant GNNs have seen success in similar areas, benefiting from the greater flexibility of their representations (Jing et al., 2020) (Satorras et al., 2021) (Du et al., 2022) (Simeon & De Fabritiis, 2024). Cartesian equivariant GNNs have also seen recent innovation in Cartesian equivariant transformer layers (Frank et al., 2022). A third significant family of geometric GNNs is spherical equivariant GNNs, which use spherical tensors rather than Cartesian tensors. As as a result, spherical equivariant GNNs behave more naturally under rotations and avail themselves of many results of the representation theory of $\mathrm{SO}(3)$. They have shown dexterity in tasks in geometry, physics, and chemistry (Thomas et al., 2018); modeled dynamic molecular systems (Anderson et al., 2019); enabled $\mathrm{SO}(3)$- and $\mathrm{SE}(3)$-equivariant transformer layers (Fuchs et al., 2020) (Liao & Smidt, 2022); accuarately and efficiently calculated interatomic potentials (Batzner et al., 2022) (Batatia et al., 2022) (Musaelian et al., 2023); enabled $E(3)$-equivariant fluid mechanical modeling  (Toshev et al., 2023); and improved efficiency by reducing certain convolution computations in $\mathrm{SO}(3)$ to equivalent ones in $\mathrm{SO}(2)$ (Passaro & Zitnick, 2023).
@@ Appendix I Extensive Details on WANet
Theorem 6 (Perturbation Bound on Eigenvalues of a Hamiltonian System).
Consider the Hamiltonian equation for a system given by
$\mathbf{H}\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon},$
where $\mathbf{H}\in\mathbb{R}^{B\times B}$ is the Hamiltonian matrix, $\mathbf{S}\in\mathbb{R}^{B\times B}$ is the overlap matrix, $\mathbf{C}\in\mathbb{R}^{B\times k}$ is the coefficient matrix, and $\bm{\epsilon}\in\mathbb{R}^{k\times k}$ is the eigenvalue matrix. Let $\hat{\mathbf{H}}=\mathbf{H}+\Delta\mathbf{H}$ be the perturbed Hamiltonian, where $\Delta\mathbf{H}\in\mathbb{R}^{B\times B}$ is a perturbation matrix.
Assume $\mathbf{S}$ is invertible, the smallest eigenvalue $\lambda_{\min}(\mathbf{S})$ of the overlap matrix $\mathbf{S}$ scales as $O(B^{-\alpha})$ for some $\alpha>0$, and the perturbation matrix $\Delta\mathbf{H}$ has entries with variance $\sigma^{2}$ and mean $\mu$. Then the perturbation in the eigenvalues $\|\hat{\bm{\epsilon}}-\bm{\epsilon}\|$ is bounded by:
$\|\hat{\bm{\epsilon}}-\bm{\epsilon}\|\leq O(B^{\alpha+\frac{1}{2}}\sigma+B^{\alpha+1}\mu).$
@@
Consider the Hamiltonian equation for the system:
$\mathbf{H}\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon},$
where $\mathbf{H}\in\mathbb{R}^{B\times B}$ is the Hamiltonian matrix, $\mathbf{S}\in\mathbb{R}^{B\times B}$ is the overlap matrix, $\mathbf{C}\in\mathbb{R}^{B\times k}$ is the coefficient matrix, and $\bm{\epsilon}\in\mathbb{R}^{k\times k}$ is the eigenvalue matrix.
Let $\hat{\mathbf{H}}=\mathbf{H}+\Delta\mathbf{H}$ be the perturbed Hamiltonian, where $\Delta\mathbf{H}\in\mathbb{R}^{B\times B}$ is a perturbation matrix. The perturbed system is given by:
$\hat{\mathbf{H}}\mathbf{C}=\mathbf{S}\mathbf{C}\hat{\bm{\epsilon}},$
Substituting $\hat{\mathbf{H}}=\mathbf{H}+\Delta\mathbf{H}$, we obtain:
$(\mathbf{H}+\Delta\mathbf{H})\mathbf{C}=\mathbf{S}\mathbf{C}\hat{\bm{\epsilon}}.$
Expanding and using the original equation $\mathbf{H}\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon}$, we have:
$\mathbf{S}\mathbf{C}\bm{\epsilon}+\Delta\mathbf{H}\mathbf{C}=\mathbf{S}\mathbf{C}\hat{\bm{\epsilon}}.$
Rearranging terms gives:
$\mathbf{S}\mathbf{C}(\hat{\bm{\epsilon}}-\bm{\epsilon})=\Delta\mathbf{H}\mathbf{C}.$
Assuming $\mathbf{S}$ is invertible, we can write:
$\mathbf{C}(\hat{\bm{\epsilon}}-\bm{\epsilon})=\mathbf{S}^{-1}\Delta\mathbf{H}\mathbf{C}.$
Taking the spectral norm on both sides, we get:
$\|\hat{\bm{\epsilon}}-\bm{\epsilon}\|\leq\|\mathbf{S}^{-1}\|_{2}\|\Delta\mathbf{H}\|_{2}.$
The spectral norm $\|\mathbf{S}^{-1}\|_{2}$ is given by the reciprocal of the smallest eigenvalue of $\mathbf{S}$:
$\|\mathbf{S}^{-1}\|_{2}=\frac{1}{\lambda_{\min}(\mathbf{S})}.$
For large $B$, the smallest eigenvalue $\lambda_{\min}(\mathbf{S})$ of the overlap matrix $\mathbf{S}$ may decrease significantly. If $\lambda_{\min}(\mathbf{S})=O(B^{-\alpha})$ for some $\alpha>0$, then:
$\|\mathbf{S}^{-1}\|_{2}=O(B^{\alpha}).$
Assuming the perturbation matrix $\Delta\mathbf{H}$ has entries with variance $\sigma^{2}$ and mean $\mu$, the spectral norm $\|\Delta\mathbf{H}\|_{2}$ scales as:
$\|\Delta\mathbf{H}\|_{2}=O(\sigma\sqrt{B}+\mu B).$
Combining these results, we have:
$\|\hat{\bm{\epsilon}}-\bm{\epsilon}\|\leq\frac{O(\sigma\sqrt{B}+\mu B)}{O(B^{-\alpha})}=O(B^{\alpha+\frac{1}{2}}\sigma+B^{\alpha+1}\mu).$
Therefore, the perturbation $\|\hat{\bm{\epsilon}}-\bm{\epsilon}\|$ grows with the dimension $B$ as $O(B^{\alpha+\frac{1}{2}}\sigma+B^{\alpha+1}\mu)$, indicating increased sensitivity to perturbations in high-dimensional systems when the overlap matrix $\mathbf{S}$ is not well-conditioned.
@@
Appendix M Improved simultaneous reduction of a matrix pair
Algorithm 2 Improved simultaneous reduction of a matrix pair $(\mathbf{H}^{\star},\mathbf{S})$
1:Ground-truth Hamiltonian matrix $\mathbf{H}^{\star}$ and overlap matrix $\mathbf{S}$
2:Diagonal matrix $\bm{\epsilon}^{\star}$ and matrix $\mathbf{C}^{\star}$ such that $(\mathbf{C}^{\star})^{\top}\mathbf{S}\mathbf{C}^{\star}=\mathbf{I}$ and $(\mathbf{C}^{\star})^{\top}\mathbf{H}^{\star}\mathbf{C}^{\star}=\bm{\epsilon}^{\star}$
3:Compute the eigen decomposition $\mathbf{V}^{\top}\mathbf{S}\mathbf{V}=\bm{\Sigma}$, where $\bm{\Sigma}$ is diagonal.
4:Define $\mathbf{G}=\mathbf{V}\bm{\Sigma}^{-1/2}$.
5:Compute $\mathbf{M}^{\star}=\mathbf{G}^{-1}\mathbf{H}^{\star}\mathbf{G}^{-\top}$.
6:Apply the symmetric QR algorithm to find the Schur form $(\mathbf{Q}^{\star})^{\top}\mathbf{M}^{\star}\mathbf{Q}^{\star}=\bm{\epsilon}^{\star}$.
7:Compute $\mathbf{C}^{\star}=\mathbf{G}^{-\top}\mathbf{Q}^{\star}$.

### v2

@@ 1 Introduction
Consequently, this computational intensity underscores the critical need to develop more efficient methodologies to determining the Hamiltonian without relying on SCF iterations. This challenge has catalyzed interest in leveraging deep learning to predict Hamiltonians directly from atomic configurations while adhering to the inherent symmetries of molecular systems (Unke et al., 2021a; Kochkov et al., 2021; Yu et al., 2023b; Li et al., 2022; Schütt et al., 2019; Gastegger et al., 2020; Hermann et al., 2020; Yin et al., 2024; Zhang et al., 2024; Luo et al., 2025). Notably, the Hamiltonian matrix is subject to unitary transformations under molecular rotations due to its spherical harmonics component. To tackle these transformations, researchers have developed SE(3)-equivariant neural networks such as PhisNet (Unke et al., 2021b) and QHNet (Yu et al., 2023b). These networks employ high-order spherical tensors and the Clebsch-Gordon tensor product to construct the predicted Hamiltonians.
@@ 2 Background
For computational efficiency, these orbitals are represented using a basis set $\{\phi_{\alpha}(\mathbf{r})\}_{\alpha=1}^{B}$ depending on the molecular geometry, $B$ denotes the number of basis and varies for different basis sets. The expansion coefficients of the orbitals are organized in a matrix $\mathbf{C}\in\mathbb{R}^{B\times N}$, allowing each orbital $\psi_{i}(\mathbf{r})$ to be expressed as: $\psi_{i}(\mathbf{r})=\sum_{\alpha=1}^{B}C_{\alpha i}\phi_{\alpha}(\mathbf{r}).$ DFT seeks the minimal electronic energy by solving for the optimal coefficient matrix $\mathbf{C}$ via the Kohn-Sham equations, represented as $\mathbf{H}(\mathbf{C})\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon},$ where $\mathbf{H}(\mathbf{C})\in\mathbb{R}^{B\times B}$ denotes the Hamiltonian matrix. Notice that $\mathbf{H}$ is a function of $\mathbf{C}$, and can be computed using density fitting with a complexity of $O(B^{3})$. $\mathbf{S}\in\mathbb{R}^{B\times B}$ is the overlap matrix with elements $S_{\alpha\beta}:=\int\phi^{\dagger}_{\alpha}(\mathbf{r})\phi_{\beta}(\mathbf{r})\mathrm{d}\mathbf{r}$, where $\phi^{\dagger}$ denotes the complex conjugate of $\phi$. $\bm{\epsilon}$ represents a diagonal matrix containing the orbital energies. This forms a generalized eigenvalue problem, where $\mathbf{C}$ comprises the eigenvectors and the diagonal elements of $\bm{\epsilon}$ are the eigenvalues. However, the solution to this problem is complicated by the interdependence between $\mathbf{H}(\mathbf{C})$ and $\mathbf{C}$. To resolve this, traditional DFT employs the self-consistent field (SCF) method, an iterative process that refines the coefficients $\mathbf{C}^{(k)}$ through successive approximations of the Hamiltonian matrix $\mathbf{H}^{(k)}$. Each iteration commences by computing $\mathbf{H}^{(k)}$ leveraging $\mathbf{C}^{(k-1)}$ and solves a new eigenvalue problem:
@@ 3 Wavefunction-Alignment Loss
Assume $\lambda_{\text{min}}(S)$ of the overlap matrix $S$ scales as: $\lambda_{\text{min}}(S)=c+\frac{A}{1+\left(\dfrac{B}{N_{0}}\right)^{\alpha}},$ where $c>0$, $A>0$, and $\alpha>0$ describe the saturation behavior. Then, for a perturbation matrix $\Delta H$, under the first-order perturbation approximation, the eigenvalue perturbation is bounded by:
$\left|\lambda_{i}(\hat{\mathbf{H}},\mathbf{S})-\lambda_{i}(\mathbf{H},\mathbf{S})\right|\leq O\left(\sigma B^{1/2}+B|\mu|\right),$
Proof.
The proof is delegated to Appendix J.2. ∎
@@
In light of this, we propose a new loss function: the Wavefunction Alignment Loss. It is designed to preserve the integrity of the eigenstructure related to molecular orbitals. Let $\hat{\bm{\epsilon}}$ and $\bm{\epsilon}^{\star}$ represent the eigenvalues (orbital energies) of $\hat{\mathbf{H}}$ and $\mathbf{H}^{\star}$, respectively, and $\hat{\mathbf{C}}$ and $\mathbf{C}^{\star}$ denote the corresponding eigenvectors (basis coefficients). We define a primary form of the loss function by directly applying the Frobenius norm to the eigenvalues $\mathcal{L}_{\mathrm{align}}=\frac{1}{n}\sum_{i=1}^{n}\|\hat{\bm{\epsilon}}_{{}_{(i)}}-\bm{\epsilon}_{{}_{(i)}}^{\star}\|^{2}_{\mathrm{F}},$ where $\hat{\bm{\epsilon}}_{{}_{(i)}}$ and $\bm{\epsilon}_{{}_{(i)}}^{\star}$ are derived from solving the generalized eigenvalue problems: $\hat{\mathbf{H}}\hat{\mathbf{C}}=\mathbf{S}\hat{\mathbf{C}}\hat{\bm{\epsilon}}$ and $\mathbf{H}^{\star}\mathbf{C}^{\star}=\mathbf{S}\mathbf{C}^{\star}\bm{\epsilon}^{\star}$, respectively.
@@
When the overlap matrix $\mathbf{S}$ is ill-conditioned, the eigenvalues $\bm{\epsilon}^{\star}$ computed by Algorithm 1 can suffer from significant roundoff errors. To mitigate this, we modify the algorithm by replacing the Cholesky decomposition of $\mathbf{S}$ with its eigen (Schur) decomposition $\mathbf{V}^{\top}\mathbf{S}\mathbf{V}=\mathbf{\Sigma}$, where $\mathbf{V}$ is the matrix of eigenvectors and $\mathbf{\Sigma}$ is the diagonal matrix of eigenvalues. We then substitute $\mathbf{G}$ with $\mathbf{V\Sigma}^{-1/2}$. This modification effectively reorders the entries of $\mathbf{M}^{\star}$, placing larger values towards the upper left-hand corner, thereby enhancing the precision in computing smaller eigenvalues (Golub & Van Loan, 2013; Wilkinson, 1988).
@@
[…]
@@ 6 Conclusion and Limitations
Acknowledgment
Tha authors would like to thank Gaoyuan Wang from Yale University for the discussion on the scaling of the eigenvalues. This work is supported by ALW professorship funds to Mark Gerstein.
@@ Appendix B Additional Related Work
Equivariant Graph Neural Networks (EGNNs)  It is often desired that machine learning (ML) models exhibit equivariance to rotations, translations, or reflections, which guarantee that they respect certain physical symmetries. Foundational work introduced group equivariant convolutional neural networks (ECNNs) (Cohen & Welling, 2016), whose importance was underscored by a proof that equivariance and convolutional structure are equivalent given certain ordinary constraints (Kondor & Trivedi, 2018). One appealing way to implement convolution is with geometric graph neural networks (geometric GNNs), which apply naturally to atomic systems by encoding them as graphs embedded in $\mathbb{R}^{3}$. Two important families of geometric GNNs are invariant GNNs and Cartesian equivariant GNNs. Over the past several years, invariant GNNs have achieved state-of-the-art results in predicting properties of molecules, crystals, and other materials  (Schütt et al., 2018) (Sanyal et al., 2018) (Chen et al., 2019) (Gasteiger et al., 2020) (Liu et al., 2022) (Gasteiger et al., 2021) (Wang et al., 2022), as well as in predicting the folding structure of proteins (Jumper et al., 2021). Cartesian equivariant GNNs have seen success in similar areas, benefiting from the greater flexibility of their representations (Jing et al., 2020) (Satorras et al., 2021) (Du et al., 2022) (Simeon & De Fabritiis, 2024). Cartesian equivariant GNNs have also seen recent innovation in Cartesian equivariant transformer layers (Frank et al., 2022). A third significant family of geometric GNNs is spherical equivariant GNNs, which use spherical tensors rather than Cartesian tensors. As as a result, spherical equivariant GNNs behave more naturally under rotations and avail themselves of many results of the representation theory of $\mathrm{SO}(3)$. They have shown dexterity in tasks in geometry, physics, and chemistry (Thomas et al., 2018); modeled dynamic molecular systems (Anderson et al., 2019); enabled $\mathrm{SO}(3)$- and $\mathrm{SE}(3)$-equivariant transformer layers (Fuchs et al., 2020) (Liao & Smidt, 2022); accuarately and efficiently calculated interatomic potentials (Batzner et al., 2022) (Batatia et al., 2022) (Musaelian et al., 2023); enabled $E(3)$-equivariant fluid mechanical modeling  (Toshev et al., 2023); and improved efficiency by reducing certain convolution computations in $\mathrm{SO}(3)$ to equivalent ones in $\mathrm{SO}(2)$ (Passaro & Zitnick, 2023).  Li et al. (2025) propose to use $\mathrm{Wiger}$-$6j$ Convolution to shift computation from edges to nodes.
@@ Appendix I Extensive Details on WANet
Theorem 6 (Perturbation Sensitivity Scaling).
Consider the generalized eigenvalue problem $\mathbf{H}\mathbf{C}=\mathbf{S}\mathbf{C}\bm{\epsilon}$, where $\mathbf{H}\in\mathbb{R}^{B\times B}$ is a symmetric Hamiltonian matrix, $\mathbf{S}\in\mathbb{R}^{B\times B}$ is a positive definite overlap matrix, $\mathbf{C}\in\mathbb{R}^{B\times k}$ is the eigenvector matrix, and $\bm{\epsilon}\in\mathbb{R}^{k\times k}$ is a diagonal matrix of eigenvalues $\lambda_{i}(\mathbf{H},\mathbf{S})$. Let $\hat{\mathbf{H}}=\mathbf{H}+\Delta\mathbf{H}$ be the perturbed Hamiltonian, with $\Delta\mathbf{H}\in\mathbb{R}^{B\times B}$ a perturbation matrix whose entries have mean $\mu$ and variance $\sigma^{2}$. Suppose the smallest eigenvalue of $\mathbf{S}$ satisfies:
$\lambda_{\min}(\mathbf{S})=c+\frac{A}{1+\left(\frac{B}{N_{0}}\right)^{\alpha}},$
where $c>0$, $A>0$, $\alpha>0$, and $N_{0}>0$ are fixed constants. Then, for each eigenvalue, there exists a constant $K>0$ such that:
$\left|\lambda_{i}(\hat{\mathbf{H}},\mathbf{S})-\lambda_{i}(\mathbf{H},\mathbf{S})\right|\leq K\left(\sigma B^{1/2}+B|\mu|\right),$
for all sufficiently large $B$.
@@
Consider the unperturbed generalized eigenvalue problem $\mathbf{H}\mathbf{c}_{i}=\lambda_{i}\mathbf{S}\mathbf{c}_{i}$, where $\mathbf{H}$ and $\mathbf{S}$ are symmetric, $\mathbf{S}$ is positive definite, and $\mathbf{c}_{i}^{T}\mathbf{S}\mathbf{c}_{i}=1$. The perturbed system is $\hat{\mathbf{H}}\hat{\mathbf{c}}_{i}=\hat{\lambda}_{i}\mathbf{S}\hat{\mathbf{c}}_{i}$, with $\hat{\mathbf{H}}=\mathbf{H}+\Delta\mathbf{H}$, $\hat{\mathbf{c}}_{i}=\mathbf{c}_{i}+\delta\mathbf{c}_{i}$, and $\hat{\lambda}_{i}=\lambda_{i}+\delta\lambda_{i}$. Substituting and expanding yields:
$(\mathbf{H}+\Delta\mathbf{H})(\mathbf{c}_{i}+\delta\mathbf{c}_{i})=(\lambda_{i}+\delta\lambda_{i})\mathbf{S}(\mathbf{c}_{i}+\delta\mathbf{c}_{i}).$
Subtracting the unperturbed equation and projecting onto $\mathbf{c}_{i}^{T}$, noting $\mathbf{c}_{i}^{T}(\mathbf{H}-\lambda_{i}\mathbf{S})=0$, we obtain:
$\mathbf{c}_{i}^{T}\Delta\mathbf{H}\mathbf{c}_{i}+\mathbf{c}_{i}^{T}\Delta\mathbf{H}\delta\mathbf{c}_{i}=\delta\lambda_{i}+\delta\lambda_{i}\mathbf{c}_{i}^{T}\mathbf{S}\delta\mathbf{c}_{i}.$
Solving for the perturbation:
$\delta\lambda_{i}=\frac{\mathbf{c}_{i}^{T}\Delta\mathbf{H}\mathbf{c}_{i}+\mathbf{c}_{i}^{T}\Delta\mathbf{H}\delta\mathbf{c}_{i}}{1+\mathbf{c}_{i}^{T}\mathbf{S}\delta\mathbf{c}_{i}}.$
For small $\Delta\mathbf{H}$, $\delta\mathbf{c}_{i}$ is $O(\|\Delta\mathbf{H}\|_{2})$, so $\mathbf{c}_{i}^{T}\Delta\mathbf{H}\delta\mathbf{c}_{i}$ and $\mathbf{c}_{i}^{T}\mathbf{S}\delta\mathbf{c}_{i}$ are higher-order terms. Here, we only consider the first-order term arises from evaluating the perturbation with the unperturbed eigenvector. Higher-order terms account for eigenvector shifts. We claim that analyzing the higher-order terms gives rise to unnecessary complication and does not alter the nature of catastrophic scaling. Specifically:
$\delta\lambda_{i}\approx\mathbf{c}_{i}^{T}\Delta\mathbf{H}\mathbf{c}_{i}.$
Bounding this term:
$|\delta\lambda_{i}|\leq|\mathbf{c}_{i}^{T}\Delta\mathbf{H}\mathbf{c}_{i}|\leq\|\mathbf{c}_{i}\|_{2}^{2}\|\Delta\mathbf{H}\|_{2}.$
Since $\mathbf{c}_{i}^{T}\mathbf{S}\mathbf{c}_{i}=1$ and $\mathbf{S}\succeq\lambda_{\min}(\mathbf{S})\mathbf{I}$, we have:
$\|\mathbf{c}_{i}\|_{2}^{2}\leq\frac{1}{\lambda_{\min}(\mathbf{S})}=\frac{1}{c+\frac{A}{1+\left(\frac{B}{N_{0}}\right)^{\alpha}}}=\frac{1+\left(\frac{B}{N_{0}}\right)^{\alpha}}{c\left[1+\left(\frac{B}{N_{0}}\right)^{\alpha}\right]+A}.$
For $\Delta\mathbf{H}$ as a random matrix, apply a spectral norm bound (by Bai-Yin law):
$\|\Delta\mathbf{H}\|_{2}\leq C_{1}(\sigma B^{1/2}+B|\mu|),$
with constant $C_{1}$, holding with high probability. Combining:
$|\delta\lambda_{i}|\leq\frac{1+\left(\frac{B}{N_{0}}\right)^{\alpha}}{c\left[1+\left(\frac{B}{N_{0}}\right)^{\alpha}\right]+A}C_{1}(\sigma B^{1/2}+B|\mu|).$
Thus, the leading-order bound is:
$\left|\lambda_{i}(\hat{\mathbf{H}},\mathbf{S})-\lambda_{i}(\mathbf{H},\mathbf{S})\right|=O\left(\frac{1+\left(\frac{B}{N_{0}}\right)^{\alpha}}{c\left[1+\left(\frac{B}{N_{0}}\right)^{\alpha}\right]+A}\left(\sigma B^{1/2}+B|\mu|\right)\right)=O\left(\sigma B^{1/2}+B|\mu|\right).$
Remark 1.
i y If higher-order terms like $\mathbf{c}_{i}^{T}\Delta\mathbf{H}\delta\mathbf{c}_{i}$ and the denominator correction are included, the perturbation includes an additional term of order $O\left(\left(\frac{1+\left(\frac{B}{N_{0}}\right)^{\alpha}}{c\left[1+\left(\frac{B}{N_{0}}\right)^{\alpha}\right]+A}\right)^{2}(\sigma B^{1/2}+B|\mu|)^{2}\right)$, reflecting the quadratic contribution of the perturbation strength.